OpenStudy (anonymous):

The equation y=-1777x+27,153 can be used to predict the number y of gun deaths in the us. x years after 2000, that is, x=o corresponds to 2000, x=2 corresponds to 2002, x=4 corresponds to 2004, and so on. Predict the number of gun deaths in 2004 and 2008. In what year will the number of gun deaths be 9383? gun deaths for 2004? gun deaths for 2008? predicted number of gun deaths will be 9383 in the year? I really need help on this one

OpenStudy (anonymous):

for the first one replace x by 4

OpenStudy (anonymous):

and use a calculator to find \[-1777\times 4+27153\]

OpenStudy (anonymous):

i get 20045

OpenStudy (anonymous):

for the next one replace x by 8 and for the last one set \[9383=-17777x+27153\] and solve for x

OpenStudy (anonymous):

So basically this equation tells you how many deaths are predicted to happen. If you think about it, when it is year 2000, x will be 0, thus the -1777x will be 0 as well, and y will be the +27,153. Now, if you you want 2001, that is x=1 so you do 27,153 - 1777*1= 27153-1777. now, for 2002, x = 2 you would get y= 27,153 - 1777*2 This is actually: 27,153 - 1777 - 1777. so each year 1777 less people die from guns each year. That's how it works. Just multiply -1777 by x and add the 27153 to get y. now, if you are given a y, you must subtract 27,153 so you get y-27153 = -1777x then you divide by -1777 in order to get x: x= (y-27153)/(-1777)

OpenStudy (anonymous):

oops i meant \[9383=-1777x+27153\]

OpenStudy (anonymous):

i get \[-17770=-1777x\] \[x=\frac{-17770}{-1777}=10\] so year would be 2010